OpenStudy (anonymous):

Amanda earned a score of 940 on a national achievement test that was normally distributed. The mean test score was 850 with a standard deviation of 100. What proportion of students had a higher score than Amanda? Use your z table. Question 6 options: 0.82 0.10 0.18 0.32

OpenStudy (anonymous):

ageta

OpenStudy (anonymous):

@ageta

OpenStudy (ageta):

tnxs

OpenStudy (boldjon):

To solve this question, you'll need to use a table of the standardized Normal distribution. Fortunately, you can just use a graphing calculator, or you can find a Normal distribution calculator on the internet. It'll ask for the three facts from your question. (If the calculator also asks for cumulative

OpenStudy (boldjon):

comulative probability

OpenStudy (boldjon):

Standard score 940 Mean 850 Standard deviation 100

OpenStudy (boldjon):

So, the answer will be the cumulative probability, which will be a number with a few decimals. However, this represents the students who had a LOWER score than Vivian. You'll need to subtract this from 1 in order to get the opposite number, the proportion of students who had a higher score than Vivian. Be sure to multiply the result by 100 to get a percentage. That's it.

OpenStudy (ageta):

σ = (940 - 850) / 100 = 0.90.

OpenStudy (boldjon):

he's wrong

OpenStudy (boldjon):

In math notation, we've done this: z = (X - μ) / σ = (940 - 850) / 100 = 0.90 where z is the z-score X is Vivian's score (940) µ is the mean (850) σ is the standard deviation (100) As you may know, in a normal distribution it's expected that about 68% of all observations will fall within 1 standard deviation of the mean, 95% will fall within 2 standard deviations, and 99% will fall within 3 standard deviations. You now have all the info you need to do a rough calculation of the cumulative probability. Here's how to put it together: The upper half of the 68% (the scores between the mean and the first standard deviation) = 34% Vivian's 90% of the 34% = 30.6% All the test scores lower than the mean = 0.500 (this is 1/2 of all test scores) + All the test scores between the mean and Vivian's score = 0.306 (see above) = 0.806, which is not the real cumulative probability, but an approximation

OpenStudy (boldjon):

0.086

OpenStudy (boldjon):

:3

OpenStudy (anonymous):

wait what

OpenStudy (anonymous):

so which one is it?

OpenStudy (anonymous):

yeah boldjon is right 0.086

OpenStudy (anonymous):

so its a

OpenStudy (anonymous):

yeah i guess

OpenStudy (anonymous):

ok preciate it